You have a parallel-plate capacitor A spark will jimp if the magnitude of the E field exceeds 3.0 x 106 V/m when air separates the plates. What is the closest distance the plates can be placed to each other without sparking? Express your answer with the appropriate units. dmin = Value Units The value should be entered as a single number or in scientific notation. No credit lost. Try again.
Answer #1

E = V/d

d = V/E = (V)/(3*10^6)

for a potential difference of V volts, the separation required is (V)/(3*10^6) m
V is not provided in the question, by entering the value of V, the distance d can be calculated.